OpenStudy (anonymous):

Lim x-> 0 2x(cot3x)(sec4x) I got to 2x(cos3x/sin3x)(1/sin4x) But I'm not sure if.. cos x/x =1 I know sin x/x =1 and x/ sin x =1 could someone help me with this problem?

OpenStudy (nikita2):

cos(x)/x -> inf as x->0

OpenStudy (anonymous):

hmm.. okay.. can someone help me solve the problem?

OpenStudy (anonymous):

did you rewrite every thing in terms of sines and cosines?

OpenStudy (anonymous):

yes ^^^ it's up there

OpenStudy (anonymous):

2x(cos3x/sin3x)(1/sin4x)

OpenStudy (anonymous):

oops mistake secant is the reciprocal of COSINE

OpenStudy (anonymous):

you should get \[\frac{\cos(3x)}{\cos(4x)}\times \frac{2x}{\sin(3x)}\]

OpenStudy (anonymous):

why would I get that 2 up there?

OpenStudy (anonymous):

sec = 1/x.. right? x is cosine

OpenStudy (anonymous):

it is the \(2x\) in your question

OpenStudy (anonymous):

\(x\) is not cosine. cosine is cosine, \(x\) is \(x\)

OpenStudy (anonymous):

yes,, but why would it only go there?? ?

OpenStudy (anonymous):

you mean instead of where?

OpenStudy (anonymous):

2x(cos3x/sin3x)(1/sin4x) it's strange that it skipped everything else and just went over sin??

OpenStudy (anonymous):

or cos** I made a mistake there

OpenStudy (anonymous):

you have \[2x\cot(3x)\sec(4x)=2x\frac{\cos(3x)}{\sin(3x)}\times \frac{1}{\cos(4x)}\] \[=\frac{2x\cos(3x)}{\sin(3x)\cos(4x)}\] \[=\frac{\cos(3x)}{\cos(4x)}\frac{2x}{\sin(3x)}\]

OpenStudy (anonymous):

i just arranged it so i can take limits i know \(\cos(0)=1\) so we can ignore the first part, and take \[\lim_{x\to 0}\frac{2x}{\sin(3x)}=\frac{2}{3}\]

OpenStudy (anonymous):

is cos(0) always 1?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so cosx = 1?

OpenStudy (anonymous):

cosine of 0 is 1, cosine of x depends on x

OpenStudy (anonymous):

OH okay but in a limit where x-> 0, this will be true?

OpenStudy (anonymous):

cosine is a function right? so for example if i have \(f(x)=x^2\) then \(f(1)=1\) but \(f\) of some other number is different

OpenStudy (anonymous):

what would at lim x-> 0 x/cos x-1 be?

OpenStudy (anonymous):

cosine is a continuous function, so the limit as x goes to 0 of cosine is \(\cos(0)=1\)

OpenStudy (anonymous):

btw you just made be from ??? to OH right now. thank you

OpenStudy (anonymous):

that is a different story if you replace \(x\) by \(0\) in that case you get \[\frac{0}{0}\]so you have to do more work

OpenStudy (anonymous):

right, okay thank you so much!

OpenStudy (anonymous):

yw, by the way the answer to \[\lim_{x\to 0}\frac{x}{\cos(x)-1}\] is it does not exist

OpenStudy (anonymous):

oh okay, thank you I was just wondering what would happen there !